OpenStudy (anonymous):

find the domain: h(x)= 2x/ x^2-4x medal and fan will be given

OpenStudy (anonymous):

@campbell_st @ganeshie8 pls help :)

OpenStudy (anonymous):

translate the question as " solve for \(x\) \(x^2-4x=0\)"

OpenStudy (anonymous):

(x+2)(x-2)=0 @satellite73

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

\[x^2-4x=0\\ x(x-4)=0\]

OpenStudy (anonymous):

yes i corrected myself.... how do i do the line and the line notation @satellite73